# How to Learn in a Noisy World? Self-Correcting the Real-World Data Noise on Machine Translation

### v1

@@ 4 Noise Self-Correction
To overcome the limitations of truncation methods in Section 3.3, we go a step further and propose the self-correction method to gradually increase the trust of model prediction distributions to correct the ground-truth data during training. Overall, our method keeps the supervision signals from the training data to avoid clean training information loss and also progressively trusts a dynamic low entropy state of model prediction distribution to revise the data. Our work is in line with the label correction in computer vision (Wang et al., 2022; Lu and He, 2022).
@@
In this way, the new target not only keeps the original supervision signal from the training data but also the model’s prediction. denotes a weighting factor to determine how much to trust the model prediction.
@@
We set correlated with a learning time function of training iteration and model entropy :
For , the model trusts its own prediction more with a more confident prediction, i.e., lower entropy. For , it allows the model to progressively trust its self-knowledge as training goes. We use a schedule (Bengio et al., 2015) to increase as a function of the training iteration and is the number of total iterations.
where and are hyper-parameters.111 We set as and to control the range of .
In general, at the beginning of training, the model is not well-trained, and a small controls the model to rely more on the ground-truth data than its own prediction. As training progresses, the increasing allows the model to trust more in its own reliable prediction.
Sharpen the Model Prediction.
To overcome the overly uncertain model prediction on learning towards the new target in Equation 2, we sharpen the model prediction distribution by controlling the softmax temperature in . We control in a fixed way and a dynamic way to vary it inversely with . Therefore, gradually decreases as training time goes: a higher at the early stage of training can prevent the model from converging and a smaller in the later stage makes the model more confident in its output. In Section 5, we compare the performance of both fixed222We use fixed followed by Wang et al. (2022). and dynamic to self-correct the data noise, and also show the impact of varying in Section 5.2.3.
Training.
After acquiring a new target , derived from both the true data and peakier model prediction distributions, we will obtain a new training objective based on maximum likelihood estimation (MLE). The following loss function is minimized for every training token over the training corpus :
@@
| dynamic | 33.5* | 32.3* | 31.4* | 33.3* | 32.5* | 30.6* | 33.5* | 31.9* | 30.4* | |
@@
| dynamic | 22.3 | 20.3* | 20.2* | 22.1* | 19.6* | 16.2* | 22.3* | 21.9* | 19.6* | |
5 Experiments
In this section, we investigate the effectiveness of our self-correction method for translation in two experiment settings: (1) simulated noise (Section 5.2), and (2) real-world noise (Section 5.3). For simulated noise, we conduct experiments by injecting two types of noise, raw-crawl data and simulated misaligned noise, into clean translation data. For the real-world noise, we perform experiments on web-mined datasets, i.e., ParaCrawl and CCAligned, across eight language pairs.
5.1 Comparing Systems
We compare our self-correction method with the following comparing systems. Note that all the models’ details are in line with the corresponding baselines. For pre-filter methods, the training data size is determined by the number of sentence pairs to filter. For data truncation and our self-correction method, both of them use the full training corpus.
Pre-Filtering.
We select two widely-used data filters: Laser and Comet. We rank the training sentence pairs based on the scores calculated by the filter models. For the simulated noise experiments (Section 5.2), we filter out the sentence pairs with the lowest scores before training, matching the size to the injected data noise. For the real-world noise experiments (Section 5.3), we filter out 20% of the sentence pairs with the lowest scores.
Truncation.
We compare two truncation methods: (1) loss truncation (Kang and Hashimoto, 2020), (2) error norm value (el2n) truncation (Li et al., 2024). Following (Li et al., 2024), we choose the best result among three truncation fractions {0.05, 0.1, 0.2} for both loss-fraction and el2n-fraction truncation. We select the best result among three threshold values {1.3, 1.35, 1.4} for el2n-threshold truncation.333Followed by (Li et al., 2024), the starting training iteration to truncate data is set as 1500.
@@ 5.2.1 Experiment Setup
We conduct experiments on both high- and low-resource translation tasks. We use the WMT2017 (German) DeEn news translation data as the high-resource task and EnSi (Sinhala) from OPUS444https://opus.nlpl.eu/ as the low-resource task. Following Herold et al. (2022), we inject noise by replacing a portion (10%, 30%, 50%) of the clean training corpus with simulated misalignment noise or raw crawl data. The misalignment noise is generated by Algorithm 1 from the replaced portion of the clean corpus. Each translation task randomly selects the raw crawl data from the raw Paracrawl corpus555https://paracrawl.eu/. The raw data contains a mixture of naturally occurring noise, including misaligned sentences, wrong language, grammar errors, etc, and provides a realistic testbed for noise-handling methods.
Our translation model uses the fairseq (Ott et al., 2019) implementation of the Transformer-Big architecture for the high-resource task and Transformer-Base for the low-resource task. The full training details are shown in Appendix C.
@@ 5.2.2 Results
Tables 3 and 4 show the high-resource DeEn and the low-resource EnSi translation performance trained on the corpus with misalignment noise or raw crawl data. Overall, the baseline model’s BLEU scores gradually drop with increasing levels of both simulated and real-world noise.
When applying pre-filter methods under simulated noisy settings, they can achieve modest gains for both simulated misalignment and the raw-crawl data. Compared with the upper-bound oracle results when the model trained with the remaining correct clean data, pre-filters notably fall short, especially in the highly (50%) noisy setting. This performance discrepancy underscores the challenge of pre-filters in accurately distinguishing simulated misalignment noise from clean data, aligning with our findings on the low filter accuracy for misalignment noise in Table 2.
For truncation methods, el2n truncation shows better performance gains than loss truncation, indicating the effectiveness of using the whole model prediction distribution to estimate data quality rather than the sole prediction probability of the ground-truth token. However, with an increasing level of noise, el2n truncation can only achieve minimal or even no improvement, showing its limitation in distinguishing clean and noisy tokens in a highly noisy training environment. In this case, removing tokens with larger el2n values leads to a loss of clean and important training data.
Instead, our self-correction method consistently improves both low- and high-resource translation performance under different controlled simulated noise settings and outperforms other systems in most scenarios. Even when injecting 50% of noise, e.g., misaligned-Laser, our self-correction method still improves over 2 BLEU points over the low-resource baseline, and nearly 1 BLEU points over the high-resource baseline. This shows the necessity of keeping the human reference information for the data sensitive scenarios, and the idea of “revising” instead of “removing” can achieve this goal. Moreover, instead of solely relying on flawed human reference, our method demonstrates the effectiveness of using the system-generated translation. This finding supports the hypothesis on the limited effectiveness on training models exclusively towards the reference especially when it is inferior to the model generated translation (Xu et al., 2024).
5.2.3 Analysis
The Impact of Self-Correction on Clean vs. Noisy Data.
Here, we aim to analyze whether the self-correction method revises the noisy samples and also maintains the performance of clean parallel data. We sample 1K clean and misaligned-Laser sentence pairs from the DeEn training corpus and record their BLEU scores during different training epochs from baseline and self-correction model. For the misaligned-Laser, we calculate the BLEU scores according to their parallel true references. Figure 4 (Left) shows that self-correction model can correct the misalignment noise by improving BLEU points compared to the baseline model. In the meantime, the self-correction model maintains the clean parallel data performance as the baseline model without hurting the clean training information, shown in Figure 4 (Middle). This further explains the superiority of self-correction method than pre-filter or data truncation from avoiding to discard the clean and useful training data.
The Impact of Sharpening Model Prediction.
Here, we aim to analyze the impact of sharpening model prediction distribution to correct the ground truth on translation performance. We train the self-correction models on DeEn task with 30% of different types of noise, with varying values of softmax temperature . From figure 4 (right), we show that using sharpening model prediction distribution with a smaller achieves better translation performance for all noisy settings. However, the optimal varies when training with different types of noise, and thus increases the difficulty of selecting a fixed for different scenarios. This motivates us to design a dynamic , which varies automatically in a low range of entropy state over training time. The overall performance in both Section 5.2 and Section 5.3 by using a dynamic also shows its general applicability for different noise scenarios.
| enfr | enru | entr† | enes† | enbe† | enht† | ensi | enkm | Avg. | ||
| Misaligned Rate (%) | 10% | 19% | 44% | 22% | 10% | 35% | 62% | 18% | - | |
| Corpus Size (M) | 5M | 5M | 5M | 5M | 1.1M | 0.55M | 0.21M | 0.06M | - | |
| Baseline | 41.1 | 26.5 | 23.5 | 21.6 | 9.9 | 17.0 | 7.0 | 4.2 | 18.8 | |
@@ 5.2.3 Analysis
| dynamic | 42.3* | 26.7* | 24.2* | 22.8* | 10.5* | 17.4* | 7.8* | 5.0* | 19.6* |
@@ 5.3.1 Experiment Setup
We investigate two noisy web-crawled datasets: Paracrawl V7.1 and CCAligned V1.0. These two datasets exhibit varying levels of semantic misalignment rates across different language pairs (Kreutzer et al., 2022). For each dataset, we select four language pairs from high- to low-resource with varying levels of misaligned noise rate. For Paracrawl, the language pairs are: enfr (French), enru (Russian), ensi (Sinhala), and enkm (Khmer). For CCAligned, the language pairs are: entr (Turkish), enes (Spanish), enbe (Belarusian), and enht (Haitian). For the high-resource language pairs: enfr, enru, entr, enes, we randomly sample 5M sentence pairs as the training corpus. For medium and low-resource language pairs, we use the original corpus size. The validation and test sets for all tasks are from Flores101.666https://github.com/facebookresearch/flores We train on the Transformer-Big (Vaswani et al., 2017) architecture for all tasks.
@@ 5.3.2 Results
Table 5 shows the translation performance for two noisy web-crawled datasets, CCAligned V1.0 and Paracrawl V7.1, across eight language pairs with varying corpus size and misaligned rates.
Overall, our self-correction method consistently outperforms other alternative methods, including pre-filter and data truncation, with up to 1.2 BLEU and 2.0 ChrF++ points improvement over the baseline. In line with the findings in the simulated noisy experiment 5.2, we also show that pre-filter and truncation methods can achieve performance gains for high-resource tasks while limited in the low-resource tasks with high misalignment rate e.g., enht, ensi and enkm. Both of the pre-filter and truncation cannot help and even hurt the low-resource translation performance. This further emphasizes the importance of keeping ground-truth supervision for the real-world noisy low-resource tasks. For our self-correction approach, it shows superior performance especially for low-resource tasks with high misaligned rates, indicating its effectiveness in handling real-world data noise in different scenarios. Apart from data pruning, our method provides new insight on handling text noise by correcting particularly for low-resource tasks.
@@ 6 Conclusion
In this paper, we aim to address the data quality issues in the web-mined parallel corpus by proposing a noise-handling self-correction approach for machine translation tasks. Our approach focuses on the translation model’s training dynamics and revises the training supervision from the ground-truth data by the model’s prediction distribution.
To evaluate our approach in different noisy settings, we take a deep look at the primary noise source in web-mined parallel data, i.e., semantic misalignment. To quantitatively analyze the impact of real-world misalignment, we propose a process to simulate the misalignment controlled by semantic similarity. Under our simulated noisy training environment, we demonstrate that our self-correction method consistently improves the baselines with different levels of misalignment noise. Moreover, we also show that our self-correction method can effectively handle naturally occurring noise in the real-world noisy web-mined datasets across eight translation tasks, outperforming other alternative approaches, including pre-filter and data truncation. Our work also provides a critical finding on the effectiveness of leveraging the system-generated translation instead of solely relying on flawed reference data.
@@ 7 Limitation
First, we acknowledge that our work aims at learning with noisy training corpus, therefore, improvements on high-quality training datasets might be limited. Second, our work focuses on “training from scratch” for translation models, while future works can apply our method to more NLP tasks and settings, including the pre-training or fine-tuning stage for text generation. Third, we only focus on the primary source of noise, i.e., misaligned sentences, on machine translation. Even we show the effectiveness of our approach on the real-world raw crawl data which contains naturally occurring data noise, future work can fine-grained evaluate our approach on other types of noise e.g., wrong language or misordered sentences.
@@ Acknowledgments
This research was funded in part by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080. We would like to thank Vlad Niculae, David Stap, Sergey Troshin and Evgeniia Tokarchuk for their useful feedback and discussion.
@@ Appendix A Data Filters
Comet is a neural framework for training machine translation evaluation models that can function as metrics Rei et al. (2020). Their framework uses cross-lingual pre-trained language modeling that exploits information from both source input and target reference to predict the target translation quality. We use the reference-free wmt-20-qe-da Comet model as the data filter to score each sentence pair in the training corpus.
For Laser, it scores sentence pairs based on cross-lingual sentence embeddings. To calculate the Laser score for each sentence pair, we generate cross-lingual sentence embeddings using the pre-trained model from Artetxe and Schwenk (2018). The underlying system is trained as a multilingual translation system with a multi-layer bidirectional LSTM encoder and an LSTM decoder, without information about the input language on the encoder. The output vectors of the encoder are compressed into a single embedding of fixed length using max-pooling, which is the cross-lingual sentence embedding resulting from the Laser model. The assumption is that two sentences with the same meaning but from different languages will be mapped onto the same embedding vectors. We calculate the cosine similarity score between source and target sentence embeddings as the Laser score. The higher the Laser score, the more semantic similar between the source and target sentence.
For Bi-Cleaner, it is a tool in Python that aims at detecting noisy sentence pairs in a parallel corpus. It indicates the likelihood of a pair of sentences being mutual translations. Sentence pairs considered high-quality are scored near with 1 and considered as noisy are scored with 0. We use the multilingual model bitextor/bicleaner-ai-full-en-xx from HuggingFace777https://huggingface.co/bitextor/bicleaner-ai-full-en-xx for the pre-filter for all language tasks.

### v2

@@ 4 Noise Self-Correction
To overcome the limitations of truncation methods in Section 3.3, we propose the self-correction method to gradually increase the trust of the model’s prediction distributions to correct the supervision during training. Our method keeps the supervision signals from the training data to avoid clean training information loss and also progressively trusts a dynamic entropy state of the model’s prediction to revise the data. Our work is in line with label correction in computer vision (discussed in Appendix C.1).
@@
In this way, the new target keeps the original supervision signal from the training data and the model’s prediction. denotes a weighting factor that determines how much to trust the model prediction.
@@
We correlate with a learning time function of training iteration and model entropy :
For , the model trusts its prediction more when it has a more confident prediction, i.e., lower entropy. For , the model can trust its self-knowledge as training progresses. We use a schedule (Bengio et al., 2015) to increase as a function of the training iteration and as the number of total iterations.
where and are hyper-parameters111We choose and based on prior experiments, see Appendix C.2. .
In general, at the beginning of training, the model is not well-trained, and a small controls the model to rely more on the ground-truth data than its prediction. As training progresses, increasing allows the model to trust more in its reliable prediction.
@@
| dynamic | 33.5 | 32.3 | 31.4 | 33.3 | 32.5 | 30.6 | 33.5 | 31.9 | 30.4 | |
Sharpen the Model Prediction.
To overcome the overly uncertain model prediction on learning towards the new target in Equation 2, we sharpen the model prediction distribution by controlling the softmax temperature in . We control in a dynamic way to vary it inversely with . Therefore, gradually decreases as training time goes on: a higher at the early stage of training can prevent the model from converging and a smaller in the later stage makes the model more confident in its output.
Training.
After acquiring a new target , derived from both the ground truth and the model’s prediction, we obtain a new training objective based on maximum likelihood estimation (MLE). The following loss function is minimized for every training token over the training corpus :
5 Experiments
In this section, we investigate the effectiveness of our self-correction method for translation tasks in two experiment settings: simulated and real-world noisy settings. For the simulated noisy setting (Section 5.2), we conduct experiments by injecting two types of noise, raw-crawl data and simulated misaligned noise, into a clean translation corpus. For the real-world noisy setting (Section 5.3), we perform experiments on two noisy web-mined datasets, i.e., ParaCrawl and CCAligned, across different language pairs.
5.1 Comparing Systems
We compare our self-correction method with the following comparing systems:333Note that all the models’ details align with the corresponding baselines.
Pre-Filtering.
We select two widely used data filters: LASER and COMET. We rank the training sentence pairs based on the scores calculated by the filter models. For the simulated noise experiments (Section 5.2), we filter out the sentence pairs with the lowest scores before training, matching the size to the injected data noise. The training data size for pre-filter methods is 90%, 70%, and 50% of the full training corpus when injecting with 10%, 30%, and 50% of data noise. For the real-world noise experiments (Section 5.3), we filter out 20% of the sentence pairs with the lowest scores.
Truncation.
We compare two truncation methods: (1) loss truncation (Kang and Hashimoto, 2020), (2) error norm value (el2n) truncation (Li et al., 2024). Following (Li et al., 2024), we choose the best result among three truncation fractions {0.05, 0.1, 0.2} for both loss and el2n truncation. The starting iteration to truncate data is set as 1500.
@@ 5.1 Comparing Systems
| dynamic | 22.3 | 20.7 | 20.2 | 22.1 | 19.6 | 16.2 | 22.3 | 21.9 | 19.6 | |
@@ 5.2.1 Experiment Setup
We conduct experiments on both high- and low-resource translation tasks. We use the WMT2017 (German) DeEn news translation data as the high-resource task and EnSi (Sinhala) from OPUS444https://opus.nlpl.eu/ as the low-resource task.
Following Herold et al. (2022), we inject noise by replacing a portion (10%, 30%, 50%) of the clean training corpus with simulated misalignment noise or raw crawl data. The misalignment noise is generated by Algorithm 1 from the replaced portion of the clean corpus. The raw crawl data noise is randomly selected from the raw Paracrawl corpus555https://paracrawl.eu/. Specifically, the raw crawl data provides a realistic test bed for noise-handling methods since it contains a mixture of naturally occurring noise, including misaligned sentences, wrong language, grammar errors, etc.
@@ 5.2.2 Results
Tables 3 and 4 show the high-resource DeEn and the low-resource EnSi translation performance trained on the corpus with simulated misalignment or raw crawl data noise. Overall, both noises negatively impact translation models, as shown by the performance drop with increasing noise levels.
First, we show that the pre-filter COMET fails to filter Misaligned-LASER noise, leading to a drop in translation performance in both high-resource and low-resource scenarios. This finding aligns with Bane et al. (2022), which demonstrates that COMET is weak at detecting misaligned segments. On the other hand, the pre-filter LASER is effective in handling Misaligned-COMET noise but only achieves modest gains when dealing with raw-crawl data noise.
Second, we demonstrate the effectiveness of leveraging the model’s self-knowledge to detect data noise during training. Consistent with our findings in Section 3.3, we show that using the el2n metric yields better performance compared to using loss. However, el2n truncation still falls short in highly noisy environments (). In such cases, the noisy datasets prevent the model from acquiring accurate knowledge, leading to incorrect data removal during training.
Our self-correction method overcomes the limitations of el2n truncation by “revising” rather than “ignoring” data noise. This approach retains the ground truth supervision, preventing the loss of clean data information. This advantage is reflected in the superior performance of self-correction across low- and high-resource tasks in various noise settings. For instance, when injecting 50% Misaligned-LASER noise into the EnSi task, our self-correction method outperforms el2n truncation by BLEU points.
Overall, our findings highlight the importance of utilizing the model’s own predictions. This supports the hypothesis that training models solely on reference translations can limit performance, particularly when the reference is inferior to the model-generated translation (Xu et al., 2024).
| enfr | entr† | enes† | enbe† | ensi | ensw | enkm | Avg. | ||
| Misaligned Rate (%) | 10% | 44% | 22% | 10% | 62% | 11% | 18% | - | |
| Corpus Size (M) | 5M | 5M | 5M | 1.1M | 210K | 130K | 60K | - | |
| Baseline | 41.1* | 23.5* | 21.6* | 9.9* | 7.0* | 13.0* | 4.2* | 17.1 | |
@@
| dynamic | 42.3 | 24.2 | 22.8 | 10.5 | 7.8 | 15.1 | 5.0 | 18.2 |
5.2.3 The Sources of Improvements
The previous section shows the benefits of the self-correction method in the presence of simulated misalignment noise. To further investigate whether the improvements arise from addressing the misaligned data, we compare the differences in translation performance on clean and Misaligned-LASER data after applying the self-correction method.
Specifically, we sample K clean and Misaligned-LASER sentence pairs and report their BLEU score differences between the baseline and the self-correction model during training. For Misaligned-LASER noisy data, BLEU scores are computed using the original parallel true references. Figure 3 indicates that the effectiveness of our self-correction method primarily stems from improving the translation quality of misaligned data. Our method enhances performance on misaligned noisy data by up to 1.5 BLEU points during training, while its impact on clean data remains minimal.
@@ 5.3.1 Experiment Setup
We investigate two noisy web-crawled datasets: Paracrawl V7.1 and CCAligned V1.0. These two datasets exhibit varying semantic misalignment rates across different low- and high-resource language pairs (Kreutzer et al., 2022). For each dataset, we select language pairs with varying levels of misaligned noise rate, from high- to low-resource. Training data details for the selected language pairs are shown in Appendix D.2.2. The validation and test sets for all tasks are from Flores101666https://github.com/facebookresearch/flores. We train for all tasks on the Transformer-Big (Vaswani et al., 2017) architecture.
@@ 5.3.2 Results
Table 5 shows the translation performance for two noisy web-crawled datasets, CCAligned V1.0 and Paracrawl V7.1, across language pairs with varying corpus size and misaligned rates.
Similar to the findings in the simulated noise setting in Section 5.2, we show that pre-filters and data truncation methods are limited to low-resource tasks with varying misalignment rates, e.g., ensw, ensi, and enkm, even degrading the translation performance. These two methods handle the data noise by removing or ignoring it; however, the noisy examples might still be partially helpful for the model, especially in data-scarce scenarios.
In contrast, the self-correction method consistently outperforms other alternative methods, including pre-filters and truncation, with an overall improvement of BLEU, COMET, and ChrF++ points over the baseline. Specifically, self-correction shows superior performance in low-resource tasks, with up to BLEU and COMET points over the baseline for ensw task. This further emphasizes the effectiveness of using the model’s self-knowledge to “correct” noise in real-world web-mined datasets.
@@ 6 Conclusion
In this paper, we aim to address the data quality issue in the web-mined translation corpora. We show that the primary noise source in translation corpora, namely semantic misalignment, is hard to filter or handle by both widely used pre-filters and data truncation methods. To quantitatively analyze the impact of misalignment noise, we propose a process to simulate it controlled by semantic similarity, which reflects the partially shared meanings often found in misaligned sentence pairs from real-world web-crawled corpora.
Under our simulated misalignment noise setting, we observe increasing reliability of the model’s self-knowledge for detecting misalignments at the token level. Building on this, we propose self-correction, which focuses on the model’s training dynamics and revises the training supervision from the reference data by the model’s prediction. Comprehensive experiments demonstrate the effectiveness of our approach on both simulated and real-world web-mined translation corpora.This performance outperforms alternative methods, including pre-filters and truncation methods. Moreover, we clearly show that the gains are mainly from revising the misaligned samples while maintaining the performance on clean data. Overall, our work provides a critical finding on the effectiveness of leveraging the model’s predictions instead of solely relying on flawed reference data.
@@ 7 Limitation
First, we acknowledge the potential bias in our self-correction method, which could learn towards the noise due to its reliance on ground truth during the early training stages. However, we believe this is not a significant issue because our method consistently demonstrates robust experimental results across different noise scenarios. Future work could explore modifications to mitigate this potential bias and enhance performance in diverse settings.
Second, our work aims at learning from a noisy training corpus, which might limit improvements when using high-quality training datasets. Furthermore, the self-correction approach has shown promise in machine translation tasks, but another limitation is the unexplored potential for other natural language processing tasks, e.g., summarization or text generation. Future work should investigate the effectiveness of this approach across different downstream tasks.
@@ Acknowledgments
This research was funded in part by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080. We would like to thank Vlad Niculae, David Stap, Sergey Troshin and Evgeniia Tokarchuk for their useful suggestions. We would also like to thank the reviewers for their feedback.
@@ Appendix A Data Filters
For LASER (Artetxe and Schwenk, 2018), it scores sentence pairs based on cross-lingual sentence embeddings. To calculate the LASER score for each sentence pair, we generate cross-lingual sentence embeddings using the pre-trained LASER model777https://github.com/facebookresearch/LASER/blob/main/nllb/README.md. The underlying system is trained as a multilingual translation system with a multi-layer bidirectional LSTM encoder and an LSTM decoder without information about the input language on the encoder. The output vectors of the encoder are compressed into a single embedding of fixed length using max-pooling, which is the cross-lingual sentence embedding resulting from the LASER model. The assumption is that two sentences with the same meaning but from different languages will be mapped onto the same embedding vectors. We calculate the LASER score followed by (Chaudhary et al., 2019). The higher the LASER score, the more semantic similar between the source and target sentence.
For COMET, it is a neural framework for training machine translation evaluation models that can function as metrics Rei et al. (2020). Their framework uses cross-lingual pre-trained language modeling that exploits information from both the source input and the target reference to predict the target translation quality. We use the reference-free wmt-20-qe-da COMET model as the data filter to score each sentence pair in the training corpus.
For Bi-Cleaner, it is a tool in Python that aims at detecting noisy sentence pairs in a parallel corpus. It indicates the likelihood of a pair of sentences being mutual translations. Sentence pairs considered high-quality are scored near , and those considered noisy are scored with . We use the multilingual model bitextor/bicleaner-ai-full-en-xx from HuggingFace888https://huggingface.co/bitextor/bicleaner-ai-full-en-xx for the pre-filter for all language tasks.
For XLM-R, it is a transformer-based multilingual masked language model pre-trained on text in 100 languages. We extract the sentence embeddings from the source and target with the model from Conneau et al. (2019) and calculate their cosine similarity score as the XLM-R score.